# Achieving Linear Speedup with ProxSkip in Distributed Stochastic Optimization

### v3

@@ VII Proof of Corollary 2
$c_{0}=4(f(\bar{{\bf{x}}}^{0})-f^{*}),\ c_{1}=\frac{2L\sigma^{2}}{n},\ c_{2}=\frac{L^{2}\sigma^{2}\big{(}\chi^{3}+256\chi(2\chi^{2}+(1-p))\big{)}}{2(1-\lambda_{2})\chi^{2}},\ a_{0}=\frac{128\chi^{2}L^{2}\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}}.$ (63)
@@
$\alpha\leq\frac{1}{\underline{\alpha}}=\min\left\{\frac{1}{2L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\}=\mathcal{O}\left(\frac{1-\lambda_{2}}{\chi L}\right).$
@@
$\frac{1}{R}\sum_{r=0}^{R-1}\mathcal{E}_{r}\leq 2c_{1}^{\frac{1}{2}}\left(\frac{c_{0}}{R}\right)^{\frac{1}{2}}+2c_{2}^{\frac{1}{3}}\left(\frac{c_{0}}{R}\right)^{\frac{2}{3}}+\frac{\left(\underline{\alpha}c_{0}+a_{0}/\underline{\alpha}^{2}\right)}{R}$ (64)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\mathcal{O}\left(\sqrt{\frac{L(f(\bar{{\bf{x}}}^{0})-f^{\star})\sigma^{2}}{nT}}\right)+\mathcal{O}\left(\sqrt[3]{\frac{\chi^{3}+\chi(1-p)}{(1-\lambda_{2})\chi^{2}}}\left(\frac{L(f(\bar{{\bf{x}}}^{0})-f^{\star})\sigma}{T}\right)^{\frac{2}{3}}\right)+\mathcal{O}\left(\frac{\frac{\chi L(f(\bar{{\bf{x}}}^{0})-f^{\star})}{1-\lambda_{2}}+\varsigma_{0}^{2}}{T}\right),$
@@
VIII Proof of Theorem 4
@@
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{8n\alpha^{4}L^{3}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}(f(\bar{{\bf{x}}}^{t})-f^{\star})+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+(1-p))}{\chi^{2}}.$ (65)
@@
$\alpha\leq\min\left\{\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},\ \chi\geq\frac{288(1-p)}{1-\lambda_{2}}\Longrightarrow\tilde{\gamma}\leq\frac{1+\gamma}{2}<1.$
@@
Since $\frac{1}{1-\gamma}\leq\frac{4\chi}{1-\lambda_{2}}$ and $\frac{1}{1-\tilde{\gamma}}\leq\frac{2}{1-\gamma}$, we have
$\frac{\frac{2\chi}{1-\lambda_{2}}}{(1-\gamma)^{2}}\leq\frac{32\chi^{3}}{(1-\lambda_{2})^{3}}\text{ and }\frac{24\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}\leq\frac{96\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\gamma)^{2}}.$
@@
$\alpha\leq\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{24\chi^{3}}}\frac{1}{4L}\Rightarrow\frac{1}{2}<1-\frac{48\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}.$
@@
$+\frac{6\alpha L}{n(1-\tilde{\gamma})}\Big{(}\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{8n\alpha^{4}L^{3}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}(f(\bar{{\bf{x}}}^{t})-f^{\star})+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+(1-p))}{\chi^{2}}\Big{)}$
$=$ $\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{6\alpha L}{n(1-\tilde{\gamma})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}-\alpha\Big{(}1-\frac{48\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}\Big{)}(f(\bar{{\bf{x}}}^{t})-f^{\star})$
$+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}$
$\leq$ $\mathcal{L}_{\mathrm{c}}^{t}-\frac{\alpha}{2}(f(\bar{{\bf{x}}}^{t})-f^{\star})+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$
@@
$\mathbb{E}\!\left[\mathcal{L}_{\mathrm{c}}^{t+1}\right]\leq\mathbb{E}\!\left[\mathcal{L}_{\mathrm{c}}^{t}\right]-\frac{\alpha}{2}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$ (66)
@@
$\frac{\alpha}{2}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\mathcal{L}_{\mathrm{c}}^{0}+T\Big{(}+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}\Big{)},$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\frac{2\mathcal{L}_{\mathrm{c}}^{0}}{\alpha T}+\frac{2\alpha\sigma^{2}}{n}+\frac{24\alpha^{4}L^{3}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{24\alpha^{2}L\sigma^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$ (67)
@@
$\mathcal{L}_{\mathrm{c}}^{0}$ $=\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}+\frac{6\alpha L}{n(1-\tilde{\gamma})}\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\leq\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}+\frac{96\chi^{2}\alpha^{3}L\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}}.$ (68)
@@
$\tilde{\gamma}\leq\frac{1+\gamma}{2}<1,\ \frac{1}{1-\gamma}\leq\frac{4\chi}{1-\lambda_{2}},$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\frac{2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{\alpha T}+\frac{192\chi^{2}\alpha^{2}L\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2\alpha\sigma^{2}}{n}+\frac{1536\chi^{3}\alpha^{4}L^{3}\sigma^{2}}{n(1-\lambda_{2})^{3}}+\frac{192\alpha^{2}L\sigma^{2}\chi(2\chi^{2}+(1-p))}{(1-\lambda_{2})\chi^{2}}.$
Since $\alpha\leq\frac{1-\lambda_{2}}{32\sqrt{3}\chi L}$, we have $\frac{1536\sigma^{2}L^{3}\alpha^{4}\chi^{3}}{n(1-\lambda_{2})^{3}}\leq\frac{\alpha^{2}L\sigma^{2}\chi}{2n(1-\lambda_{2})}\leq\frac{\alpha^{2}L\sigma^{2}\chi}{2(1-\lambda_{2})}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\frac{2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{\alpha T}+\frac{192\chi^{2}\alpha^{2}L\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2\alpha\sigma^{2}}{n}+\frac{\alpha^{2}L\sigma^{2}\chi^{3}+384\alpha^{2}L\sigma^{2}\chi(2\chi^{2}+(1-p))}{2(1-\lambda_{2})\chi^{2}}.$
@@ IX Proof of Corollary 3
$\alpha\leq\frac{1}{\underline{\alpha}}=\min\left\{\frac{1}{4L},\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{24\chi^{3}}}\frac{1}{4L}\right\}=\mathcal{O}\left(\frac{1-\lambda_{2}}{\chi L}\right).$
@@
$c_{0}=2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2},\ c_{1}=\frac{2\sigma^{2}}{n},\ c_{2}=\frac{L\sigma^{2}\big{(}\chi^{3}+384\chi(2\chi^{2}+(1-p))\big{)}}{2(1-\lambda_{2})\chi^{2}},\ a_{0}=\frac{192\chi^{2}L\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}}.$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\mathcal{O}\left(\sqrt{\frac{\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}\sigma^{2}}{nT}}+\sqrt[3]{\frac{\chi^{3}+\chi(1-p)}{(1-\lambda_{2})\chi^{2}}}L^{\frac{1}{3}}\Big{(}\frac{\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}\sigma}{T}\Big{)}^{\frac{2}{3}}+\frac{\frac{\chi L\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{1-\lambda_{2}}+\varsigma_{0}^{2}}{T}\right),$

### v4

@@ XII Proof of Corollary 3
$c_{0}=4(f(\bar{{\bf{x}}}^{0})-f^{*}),\ c_{1}=\frac{2L\sigma^{2}}{n},\ c_{2}=\frac{L^{2}\sigma^{2}\big(\chi^{3}+256\chi(2\chi^{2}+p(1-p))\big)}{2p(1-\lambda_{2})\chi^{2}},\ a_{0}=\frac{128\chi^{2}L^{2}\varsigma^{2}_{0}}{p^{2}(1-\lambda_{2})^{2}}.$ (64)
@@
$\alpha\leq\frac{1}{\underline{\alpha}}=\min\left\{\frac{1}{2L},\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L},\sqrt{\frac{p(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{p^{3}(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\}=\mathcal{O}\left(\frac{p(1-\lambda_{2})}{\chi L}\right).$
@@
$\frac{1}{T}\sum_{r=0}^{T-1}\Psi_{T}\leq 2c_{1}^{\frac{1}{2}}\left(\frac{c_{0}}{T}\right)^{\frac{1}{2}}+2c_{2}^{\frac{1}{3}}\left(\frac{c_{0}}{T}\right)^{\frac{2}{3}}+\frac{\left(\underline{\alpha}c_{0}+a_{0}/\underline{\alpha}^{2}\right)}{T}$ (65)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]\leq\mathcal{O}\left(\sqrt{\frac{L(f(\bar{{\bf{x}}}^{0})-f^{\star})\sigma^{2}}{nT}}\right)+\mathcal{O}\left(\sqrt[3]{\frac{\chi^{3}+\chi p(1-p)}{p(1-\lambda_{2})\chi^{2}}}\left(\frac{L(f(\bar{{\bf{x}}}^{0})-f^{\star})\sigma}{T}\right)^{\frac{2}{3}}\right)+\mathcal{O}\left(\frac{\frac{\chi L(f(\bar{{\bf{x}}}^{0})-f^{\star})}{p(1-\lambda_{2})}+\varsigma_{0}^{2}}{T}\right),$
@@
XIII Proof of Theorem 4
@@
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{8n\alpha^{4}L^{3}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}(f(\bar{{\bf{x}}}^{t})-f^{\star})+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+p(1-p))}{\chi^{2}}.$ (66)
@@
$\alpha\leq\min\left\{\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L},\sqrt{\frac{p(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{p^{3}(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},\ \chi\geq\frac{288(1-p)}{1-\lambda_{2}}\Longrightarrow\tilde{\gamma}\leq\frac{1+\gamma}{2}<1.$
@@
Since $\frac{1}{1-\gamma}\leq\frac{4\chi}{p(1-\lambda_{2})}$ and $\frac{1}{1-\tilde{\gamma}}\leq\frac{2}{1-\gamma}$, we have
$\frac{\frac{2\chi}{p(1-\lambda_{2})}}{(1-\gamma)^{2}}\leq\frac{32\chi^{3}}{p^{3}(1-\lambda_{2})^{3}}\text{ and }\frac{24\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{(1-\tilde{\gamma})(1-\gamma)}\leq\frac{96\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{(1-\gamma)^{2}}.$
@@
$\alpha\leq\sqrt[4]{\frac{p^{3}(1-\lambda_{2})^{3}}{24\chi^{3}}}\frac{1}{4L}\Rightarrow\frac{1}{2}<1-\frac{48\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{(1-\tilde{\gamma})(1-\gamma)}.$
@@
$+\frac{6\alpha L}{n(1-\tilde{\gamma})}\Big(\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{8n\alpha^{4}L^{3}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}(f(\bar{{\bf{x}}}^{t})-f^{\star})+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+p(1-p))}{\chi^{2}}\Big)$
$=$ $\|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}+\frac{6\alpha L}{n(1-\tilde{\gamma})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}-\alpha\Big(1-\frac{48\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{(1-\tilde{\gamma})(1-\gamma)}\Big)(f(\bar{{\bf{x}}}^{t})-f^{\star})$
$+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{p(1-\lambda_{2})}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+p(1-p))}{(1-\tilde{\gamma})\chi^{2}}$
$\leq$ $\mathcal{L}_{\mathrm{c}}^{t}-\frac{\alpha}{2}(f(\bar{{\bf{x}}}^{t})-f^{\star})+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{p(1-\lambda_{2})}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+p(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$
@@
$\mathbb{E}\!\left[\mathcal{L}_{\mathrm{c}}^{t+1}\right]\leq\mathbb{E}\!\left[\mathcal{L}_{\mathrm{c}}^{t}\right]-\frac{\alpha}{2}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{p(1-\lambda_{2})}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+p(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$ (67)
@@
$\frac{\alpha}{2}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\mathcal{L}_{\mathrm{c}}^{0}+T\Big(+\frac{\alpha^{2}\sigma^{2}}{n}+\frac{12\alpha^{5}L^{3}\sigma^{2}\frac{2\chi}{p(1-\lambda_{2})}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{12\alpha^{3}L\sigma^{2}(2\chi^{2}+p(1-p))}{(1-\tilde{\gamma})\chi^{2}}\Big),$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\frac{2\mathcal{L}_{\mathrm{c}}^{0}}{\alpha T}+\frac{2\alpha\sigma^{2}}{n}+\frac{24\alpha^{4}L^{3}\sigma^{2}\frac{2\chi}{p(1-\lambda_{2})}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{24\alpha^{2}L\sigma^{2}(2\chi^{2}+p(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$ (68)
@@
$\mathcal{L}_{\mathrm{c}}^{0}$ $=\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}+\frac{6\alpha L}{n(1-\tilde{\gamma})}\|\mathcal{E}^{0}\|_{\mathrm{F}}^{2}\leq\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}+\frac{96\chi^{2}\alpha^{3}L\varsigma^{2}_{0}}{p^{2}(1-\lambda_{2})^{2}}.$ (69)
@@
$\tilde{\gamma}\leq\frac{1+\gamma}{2}<1,\ \frac{1}{1-\gamma}\leq\frac{4\chi}{p(1-\lambda_{2})},$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\frac{2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{\alpha T}+\frac{192\chi^{2}\alpha^{2}L\varsigma^{2}_{0}}{p^{2}(1-\lambda_{2})^{2}T}+\frac{2\alpha\sigma^{2}}{n}+\frac{1536\chi^{3}\alpha^{4}L^{3}\sigma^{2}}{np^{3}(1-\lambda_{2})^{3}}+\frac{192\alpha^{2}L\sigma^{2}\chi(2\chi^{2}+p(1-p))}{p(1-\lambda_{2})\chi^{2}}.$
Since $\alpha\leq\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L}$, we have $\frac{1536\sigma^{2}L^{3}\alpha^{4}\chi^{3}}{np^{3}(1-\lambda_{2})^{3}}\leq\frac{\alpha^{2}L\sigma^{2}\chi}{2np(1-\lambda_{2})}\leq\frac{\alpha^{2}L\sigma^{2}\chi}{2p(1-\lambda_{2})}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\frac{2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{\alpha T}+\frac{192\chi^{2}\alpha^{2}L\varsigma^{2}_{0}}{p^{2}(1-\lambda_{2})^{2}T}+\frac{2\alpha\sigma^{2}}{n}+\frac{\alpha^{2}L\sigma^{2}\chi^{3}+384\alpha^{2}L\sigma^{2}\chi(2\chi^{2}+p(1-p))}{2p(1-\lambda_{2})\chi^{2}}.$
@@ XIV Proof of Corollary 4
$\alpha\leq\frac{1}{\underline{\alpha}}=\min\left\{\frac{1}{4L},\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L},\sqrt{\frac{p(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L},\sqrt[4]{\frac{p^{3}(1-\lambda_{2})^{3}}{24\chi^{3}}}\frac{1}{4L}\right\}=\mathcal{O}\left(\frac{p(1-\lambda_{2})}{\chi L}\right).$
@@
$c_{0}=2\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2},\ c_{1}=\frac{2\sigma^{2}}{n},\ c_{2}=\frac{L\sigma^{2}\big(\chi^{3}+384\chi(2\chi^{2}+p(1-p))\big)}{2p(1-\lambda_{2})\chi^{2}},\ a_{0}=\frac{192\chi^{2}L\varsigma^{2}_{0}}{p^{2}(1-\lambda_{2})^{2}}.$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq$ $\mathcal{O}\left(\sqrt{\frac{\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}\sigma^{2}}{nT}}+\sqrt[3]{\frac{\chi^{3}+\chi p(1-p)}{(1-\lambda_{2})\chi^{2}}}L^{\frac{1}{3}}\Big(\frac{\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}\sigma}{T}\Big)^{\frac{2}{3}}+\frac{\frac{\chi L\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\|^{2}}{p(1-\lambda_{2})}+\varsigma_{0}^{2}}{T}\right),$